# Why Stop at One Error? Benchmarking LLMs as Data Science Code Debuggers for Multi-Hop and Multi-Bug Errors

### v2

@@ Abstract
LLMs are transforming software development, yet current code generation and code repair benchmarks mainly assess syntactic and functional correctness in simple, single-error cases. LLMs’ capabilities to autonomously find and fix runtime logical errors in complex data science code remain largely unexplored. To address this gap, we introduce DSDBench: the Data Science Debugging Benchmark, the first benchmark for systematic evaluation of LLMs on multi-hop error tracing and multi-bug detection in data science code debugging. DSDBench adapts datasets from existing data science task benchmarks, such as DABench and MatPlotBench, featuring realistic data science debugging tasks with automatically synthesized multi-hop, multi-bug code snippets. DSDBench includes 1,117 annotated samples with 741 cause-effect error pairs and runtime error messages. Evaluations of state-of-the-art LLMs on DSDBench show significant performance gaps, highlighting challenges in debugging logical runtime errors in data science code. DSDBench offers a crucial resource to evaluate and improve LLMs’ debugging and reasoning capabilities, enabling more reliable AI-assisted data science in the future. 111  DSDBench is publicly available at https://github.com/KevinCL16/DSDBench.
@@ 1 Introduction
As illustrated in Figure 1, unlike constrained programming exercises, debugging data science codebases presents unique challenges: 1) Its heavy reliance on external libraries (e.g., pandas, NumPy, scikit-learn, matplotlib) means subtle misuses or incorrect data processing steps can easily trigger downstream runtime exceptions. 2) Data scientists often work in interactive environments like Jupyter Notebooks, which lack robust debugging tools. This makes it harder to identify and fix runtime bugs, especially when multiple subtle errors, such as incorrect data transformations or misaligned indices, coexist and interact within the code, complicating the debugging process. 3) Standard debugging tools offer limited assistance in diagnosing multi-hop logical errors within complex workflows. The root cause of these errors can be distantly located from the point of error manifestation. Standard debuggers typically report the symptom (the line of error manifestation in the stack trace) rather than the root cause responsible for the program’s termination. Overall, a dedicated benchmark for rigorously assessing LLMs’ dynamic debugging of multi-hop logical errors in complex multi-bug data science code is still lacking.
Motivated by this evident gap in evaluating LLMs’ dynamic debugging skills for data science, we introduce DSDBench: the Data Science Debugging Benchmark. Distinct from prior works that primarily focus on repairing single, syntactic and static errors, where these errors are easily caught by interpreters or compilers, DSDBench is the first benchmark to systematically evaluate LLMs on: (1) Multi-Hop Error Tracing: requiring models to trace runtime errors back through multiple lines of data science code to identify the root cause; and (2) Multi-Bug Error Detection: assessing their ability to concurrently detect and reason about multiple logical errors within a single data science code snippet. Table 1 summarizes the comparisons between DSDBench and existing code debugging benchmarks.
@@ 2.2 Error Injection
To systematically introduce errors, we employ two error injection methodologies. The details regarding the error injection prompts are provided in Appendix B.
@@
One primary method is called strong LLM-based error injection, which utilizes a strong LLM, i.e., GPT-4o, to inject runtime-interrupting errors. This involves a two-stage process: 1) We identify code lines using data science libraries (numpy, scipy, matplotlib, sklearn, pandas) by instructing GPT-4o to extract relevant core library functions. 2) We inject runtime errors into these lines using GPT-4o by introducing plausible, contextually relevant runtime errors within code identified in the first step, causing programs to halt. In terms of multi-hop errors, error injection on one line could cause a runtime error to manifest on a later line due to sequential code execution.
@@
Across our three data sources, we explore weak LLM-based direct error generation using Llama-3.1-8B. We instruct Llama-3.1-8B to directly generate Python code from benchmark questions. Due to the limitations of weaker LLMs, the generated code often contains errors. In terms of multi-hop errors, in directly generated code with functions, an error in a sub-function could trigger an error reported in the main function during execution.
@@ 2.3 Error Annotation
For each buggy code snippet, we annotate three ground truths, cause_error_line, effect_error_line, and runtime error messages. Our annotation process can be divided into single-error and multi-error phases.
@@
Multi-error annotations are generated based on single-error annotations. For each question, we systematically generated all combinations of the already annotated single errors to create a pool of candidate multi-error annotations. From each question’s candidate pool, we randomly sample a subset to form our multi-error dataset.
@@ 2.5 Dataset Characteristics
Figure 2 illustrates the error type distribution in DSDBench. Figure 3 shows the data science library coverage within the dataset.
@@ 3.1 Task Definition
The benchmark is specifically designed to assess two critical dimensions of debugging proficiency: multi-hop error detection and multi-bug error detection. Multi-hop error detection evaluates the LLMs’ ability to trace errors to their root cause (cause_error_line), which may precede the interpreter’s error point (effect_error_line) by several lines of code. Multi-bug error detection assesses the LLMs’ ability to identify and explain multiple, concurrent logical errors within a single code snippet, rather than solely the initial error encountered. A further goal is to evaluate the quality of error message reproduction, specifically the ability of LLMs to accurately reproduce the error messages thrown by the Python interpreter for each identified error.
@@ 3.2 Evaluation Metrics
Specifically, we calculate cause_line_score, effect_line_score, and error_type_score as binary metrics (1 for exact match with ground truth, 0 otherwise). error_message_score is evaluated by GPT-4o on a scale in [0.0, 0.25, 0.5, 0.75, 1.0] based on the relevance and correctness of the reproduced error message compared to the ground truth error message.
@@ 4.1 Setup
The prompt used to evaluate all models are identical, including task description, buggy Python code snippet from DSDBench, and instructions to output the debugging analysis in a structured JSON format. The precise prompt template is in Appendix D. This zero-shot evaluation approach allows us to assess the inherent debugging abilities of each model. We utilized the metrics defined in Section 3.2 for quantitative assessment.
@@ 4.3 Impact of Self-Debugging
To further investigate the impact of LLM-as-debuggers on real-world coding tasks, we explored using LLM-generated debugging information in data science coding tasks as a self-refining mechanism. In this experiment, models were tasked with solving DABench-Hard either directly or by refining their initial code using self-generated debugging information.
Table 4 presents the accuracy of GPT-4o, Claude 3.5 sonnet, and Qwen2.5-72B-Instruct in various settings. Across models, Self-Refinement significantly improves accuracy compared to the Direct Solution. Furthermore, performance drops when either the cause line (No Cause) or effect line (No Effect) is removed from the debugging information. Removing only the error message (No Message) has less negative impact.

### v3

@@ Abstract
LLMs are transforming software development, yet most code benchmarks still emphasize syntactic or functional correctness in simple, single-error cases. These settings miss the core difficulty of real-world data science debugging, where runtime errors propagate across multiple lines (multi-hop) and often appear in sets (multi-bug). We introduce DSDBench: Data Science Debugging Benchmark, the first benchmark to systematically evaluate LLMs on this challenge. Unlike general debugging benchmark suites such as SWE-bench, DSDBench targets non-expert, data-centric scripting, where practitioners rely heavily on black-box libraries and write exploratory code that is error-prone and difficult to debug. Evaluations of state-of-the-art LLMs reveal large performance gaps: even frontier models that excel at code generation fail to reliably trace and resolve these errors, exposing a critical “generation versus understanding” gap. DSDBench provides a resource to drive progress toward more robust and trustworthy AI-assisted data science.  [^1] DSDBench is publicly available at https://github.com/KevinCL16/DSDBench.
@@ 1 Introduction
As illustrated in Figure 1, unlike constrained programming exercises, debugging data science codebases presents unique challenges: 1) Its heavy reliance on external libraries (e.g., pandas, NumPy, scikit-learn, matplotlib) means subtle misuses or incorrect data processing steps can easily trigger downstream runtime exceptions. 2) Data scientists often work in interactive environments like Jupyter Notebooks, which lack robust debugging tools. This makes it harder to identify and fix runtime bugs, especially when multiple subtle errors, such as incorrect data transformations or misaligned indices, coexist and interact within the code, complicating the debugging process. 3) Standard debugging tools offer limited assistance in diagnosing multi-hop logical errors within complex workflows. The root cause of these errors can be distantly located from the point of error manifestation. Standard debuggers typically report the symptom (the line of error manifestation in the stack trace) rather than the root cause responsible for the program’s termination. Linking these two is essential not only for reasoning, but also for trustworthy assistance grounded in what users actually see.
Overall, a dedicated benchmark for rigorously assessing LLMs’ dynamic debugging of multi-hop logical errors in complex multi-bug data science code is still lacking. Unlike general debugging suites such as SWE-bench that evaluate patch-level fixes in professional OSS repositories via unit tests, our setting targets non-expert, data-centric scripting where failures are often data-dependent and surface as interpreter-visible runtime errors. Crucially, we evaluate whether models can connect root causes to these interpreter-visible failure points (effect lines), a prerequisite for trustworthy explanations.
Motivated by this evident gap in evaluating LLMs’ dynamic debugging skills for data science, we introduce DSDBench: the Data Science Debugging Benchmark. Distinct from prior works that primarily focus on repairing single, syntactic and static errors, DSDBench is the first benchmark to systematically evaluate LLMs on: (1) Multi-Hop Error Tracing: requiring models to trace runtime errors back through multiple lines of data science code to identify the root cause; and (2) Multi-Bug Error Detection: assessing their ability to concurrently detect and reason about multiple logical errors within a single data science code snippet. We focus on interpreter-visible (crashing) bugs to enable scalable, deterministic annotation of cause/effect lines and messages. For clarity, we define a “hop” whenever the cause and effect lines differ, and for multi-bug items a prediction is counted correct only if all constituent bugs are identified. Table 1 summarizes the comparisons between DSDBench and existing code debugging benchmarks.
@@ 2.2 Error Injection
To systematically introduce errors, we employ two complementary methods. This dual design ensures both realism and diversity of bug types. The details of prompts are provided in Appendix B.
@@
We instruct GPT-4o to inject runtime-interrupting bugs into otherwise correct solutions. Step 1: identify lines invoking core data science libraries (numpy, scipy, matplotlib, sklearn, pandas). Step 2: modify them to introduce plausible faults such as API misuse, incorrect parameter settings, or data-shape mismatches (e.g., NaNs, inconsistent dimensions). These produce realistic logic/runtime bugs. Importantly, injection often yields multi-hop errors: the causal bug occurs earlier in the program, but the crash manifests later (e.g., a faulty imputation call causing a failure during model training).
@@
We also instruct Llama-3.1-8B to write entire solutions from scratch. Due to its weaker capability, the generated code frequently contains beginner-style mistakes (e.g., missing imports, wrong API signatures, incomplete function definitions). These naturally introduce multi-hop errors, for example when a sub-function mistake surfaces as an exception in the main function.
Rationale
Using both approaches diversifies error characteristics: strong LLM injection yields semantically rich runtime bugs, while weak LLM generation reflects novice-style mistakes. This mitigates bias from relying on a single method.
@@ 2.3 Error Annotation
For each buggy snippet, we annotate three ground truths: cause_error_line, effect_error_line, and runtime error message. We then construct both single-error and multi-error cases, requiring models to reason across causal and manifested locations.
@@
We create multi-error cases by systematically combining validated single errors into candidate pools per question, from which random subsets are sampled. An instance is marked correct only if all constituent bugs are identified. Although each bug is logically independent, runtime interactions (e.g., early crashes masking later ones) can increase difficulty, which we preserve to reflect realistic debugging.
Multi-Hop Definition
We define a “hop” as the spatial separation between the cause line (where the bug originates) and the effect line (where execution fails). The hop distance may be a few lines within the same function (no change in call stack) or span across functions (involving call stack depth). The essential characteristic is that the interpreter reports a crash at a different line than where the bug was introduced, forcing models to trace execution flow.
@@ 2.4 Human Verification and Quality Control
Bug Distribution, Realism, and Scope
We approximate non-expert data-science scripting, where code is exploratory and loosely structured. To mitigate LLM-specific biases, our injection is constrained and modular: a strong LLM first identifies API-invoking lines and then performs atomic, plausible edits (e.g., argument name/value mistakes, missing axis, shape mismatches), limiting unrealistic global rewrites and better mimicking small mistakes that propagate during execution. Weak-LLM generation complements this by introducing novice-style patterns at the program level. We presently focus on interpreter-visible (crashing) bugs to support reliable, scalable labeling; functional bugs without hard crashes are deferred to future work. Empirically, models with stronger coding/reasoning (e.g., Claude-3.5-Sonnet, DeepSeek-V3, LRMs) often outperform GPT-4o (used for injection), suggesting DSDBench rewards genuine reasoning rather than model-specific shortcuts.
@@ 2.5 Dataset Characteristics
Figure 2 summarizes the error type distribution, and Figure 3 shows the library coverage. Together they indicate broad coverage across common runtime/data-driven failures and core APIs, lending distributional realism despite synthetic injection.
@@ 3.1 Task Definition
The benchmark targets two critical dimensions: multi-hop error detection and multi-bug error detection. Multi-hop evaluates whether models can trace errors to their root cause (cause_error_line) when it differs from the interpreter’s error point (effect_error_line). We define a “hop” whenever these two lines are not identical; the distance may be within the same function (no call-stack change) or across function calls (with call-stack depth). This spatial separation forces reasoning over execution flow. Multi-bug requires identifying and explaining all concurrent logical errors in a snippet, not merely the first encountered, mirroring real debugging where missing one error leaves the program incorrect. We also evaluate error message reproduction, i.e., whether models can semantically reproduce the interpreter-thrown messages for each identified error.
@@ 3.2 Evaluation Metrics
We compute cause_line_score, effect_line_score, and error_type_score as binary exact matches (1 for exact match with ground truth, 0 otherwise). For error messages, exact string matching is too brittle: LLMs may produce semantically equivalent messages with different surface forms. We therefore adopt a semantic rubric scored by GPT-4o on a five-point Likert scale {0.0, 0.25, 0.5, 0.75, 1.0}. Example: GT: “ValueError: Input contains NaN” vs. Prediction: “ValueError: The model cannot handle NaN values in the input”—lexically different but semantically equivalent (scored $\geq 0.75$). The full rubric and additional examples are provided in Appendix D.
@@ 4.1 Setup
The evaluation prompt is identical across models and includes the task description, the buggy DSDBench Python snippet, and instructions to output a structured JSON diagnosis. We disallow external tools and execution to isolate in-context reasoning. For multi-bug items, a prediction is counted correct only if all constituent bugs are identified (no partial credit), consistent with Section 2.3. The template appears in Appendix D. Metrics follow Section 3.2.
A pilot study with agentic systems is reported in Appendix G; here we focus the main results on in-context localization without execution.
@@ 4.3 Impact of Self-Debugging
To study how DSDBench-style diagnosis helps downstream coding, we use models’ own debugging outputs as guidance to solve DABench-Hard, comparing direct solutions vs. self-refined solutions that consume (cause line, effect line, message).
Figure 4 shows that self-refinement improves accuracy across GPT-4o, Claude 3.5 Sonnet, and Qwen2.5-72B. Ablations confirm that lines carry most of the value—removing Cause/Effect hurts more than removing Message.
@@ 4.4 Detailed Analysis
Prompting Strategies
We additionally benchmark several test-time prompting strategies (Zero-shot/One-shot CoT, Self-Refine, Self-Consistency) on a single-bug subset. Self-Consistency yields the largest gains for some models, but overall performance remains far from saturation. Full setup and results appear in Appendix F.
Agentic Systems
We also ran a pilot with agentic coding systems (Cursor Agent with Claude 3.5 Sonnet, and OpenHands). Single-attempt repair boosts performance over standalone localization, while unconstrained workflows can approach 100% pass rate by broadly rewriting code. Full setup, tables, and analysis are provided in Appendix G.
@@ 6 Conclusion
Acknowledgments
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C004).
@@ Appendix E Full Evaluation Results
Appendix F Prompting Strategies and Test-Time Compute
We evaluate four test-time reasoning strategies beyond direct zero-shot prompting on the single-bug subset used for LRMs: Zero-shot CoT, One-shot CoT, Self-Refine, and Self-Consistency. All runs disallow external tools and execution to isolate in-context reasoning. See results in Table 11.
Model Prompting Cause Acc. Effect Acc. $\Delta$ Cause vs. Direct $\Delta$ Effect vs. Direct
GPT-4o Direct (Baseline) 35.4 31.2
GPT-4o Zero-shot CoT 33.3 30.2 -2.1 -1.0
GPT-4o One-shot CoT 33.3 31.3 -2.1 +0.1
GPT-4o Self-Refine 33.3 30.2 -2.1 -1.0
GPT-4o Self-Consistency 37.5 34.4 +2.1 +3.2
DeepSeek-V3 Direct (Baseline) 44.8 28.1
DeepSeek-V3 Zero-shot CoT 44.8 33.3 +0.0 +5.2
DeepSeek-V3 One-shot CoT 46.9 33.3 +2.1 +5.2
DeepSeek-V3 Self-Refine 47.9 31.3 +3.1 +3.2
DeepSeek-V3 Self-Consistency 42.7 37.5 -2.1 +9.4
Qwen2.5-72B Direct (Baseline) 26.0 26.0
Qwen2.5-72B Zero-shot CoT 37.5 31.3 +11.5 +5.3
Qwen2.5-72B One-shot CoT 37.5 25.0 +11.5 -1.0
Qwen2.5-72B Self-Refine 37.5 31.3 +11.5 +5.3
Qwen2.5-72B Self-Consistency 40.6 36.5 +14.6 +10.5
Table 11: Prompting strategies on the single-bug subset. Gains show sensitivity to test-time compute and scaffolding; however, all models remain far from saturation.
Observations. Self-Consistency is most helpful for GPT-4o; DeepSeek-V3 benefits more on effect-line localization; Qwen2.5-72B experiences large gains, reflecting stronger utility from added scaffolds. Despite improvements, the absolute accuracies remain modest, underscoring DSDBench’s difficulty.
Appendix G Agentic Systems: Pilot Evaluation
We evaluate whether agentic workflows mitigate DSDBench’s challenges using Cursor Agent (Claude 3.5 Sonnet backend) and OpenHands on the single-bug subset. We consider two settings:
  • Single-Attempt Repair (pass@1): Canonicalize each item into a one-line repair task; the agent may change exactly one line once. Success requires a clean reference run.
  • Unconstrained Agentic Workflow: The agent may iteratively invoke tools and execute code until success (no edit budget).
See results in Table 12
Model / Approach Task Metric Performance (%) Key Insight
Claude-3.5-Sonnet (Standalone) Cause Line Localization Accuracy 34.0 Baseline “understanding” is low.
Cursor Agent (Claude-3.5-Sonnet) Single-Attempt Repair pass@1 48.9 Recasting to generation yields +14.9 points.
Cursor Agent (Claude-3.5-Sonnet) Unconstrained Repair Pass Rate $\sim$100 Often succeeds via broad, non-minimal rewrites.
OpenHands (Cloud GUI) Single-Attempt Repair pass@1 18.6 Large variance across agents/backbones.
OpenHands (Cloud GUI) Unconstrained Repair Pass Rate 41.7 Iterative repair remains non-trivial.
Table 12: Agentic pilot on the single-bug subset. Backbone LLM unknown for the cloud GUI; results are indicative only.
Takeaways. (i) Single-attempt agentic repair outperforms standalone localization, indicating that reframing from diagnosis to generation aligns better with autoregressive training. (ii) Unconstrained agents can brute-force success by extensive regeneration, masking the reasoning gap DSDBench is designed to expose. This justifies our focus on line-level localization and semantic message reproduction as complementary to patch-level pass rates.